OpenStudy (anonymous):

An absolute equation x+4 <0, I am a bit lost, I thought I understood but now am confused, can someone help please......I thought no solutions am I correct?

OpenStudy (perl):

| x + 4 | < 0 ?

OpenStudy (anonymous):

yes :)

OpenStudy (perl):

youre in luck, an absolute value can never be negative, so there are no solutions

OpenStudy (anonymous):

Excellent thank you so very much :)

OpenStudy (perl):

$$ \Large |x| \geq 0 $$

OpenStudy (anonymous):

So in my justification if I say: The absolute value of x must always be greater than or equal to zero?

OpenStudy (perl):

no matter what you put inside the absolute value bars, the output will always be positive

OpenStudy (perl):

or zero

OpenStudy (perl):

$$ \LARGE | anything | \geq 0 $$

OpenStudy (anonymous):

what about \[\left| x-6 \right|\ge0\]

OpenStudy (anonymous):

I find these so confusing :(

OpenStudy (perl):

oh wait ,it must be true for all x

OpenStudy (perl):

since the absolute value of anything is positive or zero

OpenStudy (anonymous):

I was thinking that, so x is an element of all the reals?

OpenStudy (perl):

yes

OpenStudy (perl):

let me give you two general solutions

OpenStudy (perl):

$$ \large \text{assuming that} ~a \geq 0\\ \Large {|x| \leq a \iff -a \leq x \leq a \\ |x| \geq a \iff x \geq a ~~or~~ x \leq -a } $$

OpenStudy (perl):

$$ \large \text{assuming that} ~a \geq 0\\ \Large {|x| <a \iff -a<x < a \\ |x|>a \iff x >a ~~or~~ x <-a \\ \therefore | x - 6 | \geq 0 \iff x-6 \geq 0 ~~ or ~~ x-6 \leq 0 } $$

OpenStudy (perl):

now what is the solution for x >= 6 or x <=6 , that must be any real number

OpenStudy (anonymous):

yes, got it, thanks you make much more sense xoxo

OpenStudy (perl):

it is faster to just say, |x | >= 0 is true for any x

OpenStudy (anonymous):

Thank you again :)